# Music Arena: Live Evaluation for Text-to-Music

### v1

@@ 7 Limitations and Future Work
Our work has several limitations that present clear avenues for future improvement. On the frontend, our ability to precisely track which segment of the audio a user is listening to is constrained by the user interface—we track total listening duration but cannot capture seeking actions within audio clips. Our current backend system also selects pairs of TTM systems uniformly at random, rather than more principled strategies (Chi et al., 2025) that navigate tradeoffs around quality, speed, and coverage of relevant pairs. Furthermore, the scope of Music Arena is currently limited to text-to-music generation, excluding other important tasks like symbolic generation or style transfer. Finally, as a public web platform, our user base is not representative of the global population, and the long-term sustainability of providing free access to self-hosted open-weights models remains a challenge.
Our plans for future work are motivated by addressing these limitations, improving our understanding of human musical preferences, and advancing the science of live evaluation more broadly. We aim to continue refining our frontend and backend—a particular direction of interest is improving the backend pair selection algorithm to better balance tradeoffs around leaderboard fidelity and user experience. We will leverage the growing preference dataset to better understand strengths and weaknesses of specific models and perform meta-evaluation against automatic evaluation metrics. Through analysis of natural language feedback and live evaluation of controlled degradations of systems (e.g., adding latency or noise to a system), we may better understand which attributes users consider most prominently when making preference decisions. As creative workflows around music generation mature, we hope to integrate live evaluation directly into user workflows (Chi et al., 2025). Finally, we will continuously refine our evaluation methodology based on community feedback to ensure the long-term rigor and fairness of our platform.
@@ 8 Conclusion
The development of Music Arena was supported by Sony AI. We extend our sincere thanks to our commercial contacts at Riffusion, Stability AI, Google DeepMind, and Suno for productive discussions that informed the key features and policies of Music Arena.
Appendix A Complete example of a Music Arena battle
@@
In Figure˜2 we show an example of a completed battle in Music Arena. Below we include the complete JSON log from our platform for that same battle, highlighting the detailed preference information that we collect. The corresponding audio is here: https://drive.google.com/drive/folders/1UlueXyaTmef2qw5zwdctXNVgKP9bFa1R?usp=sharing

### v2

@@ 6 Limitations and Future Work
Our work has several limitations that present clear avenues for future improvement. On the frontend, our ability to precisely track which segment of the audio a user is listening to is constrained by the user interface—we track total listening duration but cannot capture seeking actions within audio clips. Our current backend system also selects pairs of TTM systems uniformly at random, rather than more principled strategies [16] that navigate tradeoffs around quality, speed, and coverage of relevant pairs. Furthermore, the scope of Music Arena is currently limited to text-to-music generation, excluding other important tasks like symbolic generation or style transfer. Finally, as a public web platform, our user base is not representative of the global population, and the long-term sustainability of providing free access to self-hosted open-weights models remains a challenge.
Our plans for future work aim to address these limitations, improve our understanding of human musical preferences, and contribute to the science of live evaluation. We aim to continue refining our frontend and backend—a particular direction of interest is improving the backend pair selection algorithm to better balance tradeoffs around leaderboard fidelity and user experience. We will leverage the growing preference dataset to better understand strengths and weaknesses of specific models and perform meta-evaluation against automatic evaluation metrics. Through analysis of natural language feedback and live evaluation of controlled degradations of systems (e.g., adding latency or noise to a system), we may better understand which attributes users consider most prominently when making preference decisions. As creative workflows mature, we hope to integrate live evaluation directly into user workflows [16]. Finally, we will continuously refine our evaluation methodology based on community feedback to ensure the long-term rigor and fairness of our platform.
@@ 7 Conclusion
The development of Music Arena was supported by Sony AI. We extend our sincere thanks to our commercial contacts at Producer.ai, Stability AI, Google DeepMind, and Suno for productive discussions that informed the key features and policies of Music Arena.
Appendix A Leaderboard structure
Table 1: The first Music Arena leaderboard results (July 28 – Aug 31, 2025), segmented by Instrumental and Vocal models. The table includes Arena Scores (with 95% confidence intervals), vote counts, generation speed (RTF), and key metadata for each model.
Instrumental Music Generation
Rank Model Arena Score 95% CI # Votes Generation Speed (RTF) Organization License Training Data Supports Lyrics Access
1 riffusion-fuzz-1-1 1250.8 +52.0 / -45.5 252 6.01 Producer.ai Closed Unspecified True Proprietary
2 magenta-rt-large 1113.6 +56.5 / -57.2 276 1.01 Google DeepMind Apache 2.0 Stock False Open weights
3 musicgen-small 928.5 +40.4 / -46.7 278 0.86 Meta CC-BY-NC 4.0 Stock False Open weights
4 sao 924.7 +45.7 / -41.5 286 2.63 Stability AI STAI Community Open False Open weights
5 sao-small 782.4 +50.9 / -62.2 292 12.79 Stability AI STAI Community Open False Open weights
Vocal Music Generation
Rank Model Arena Score 95% CI # Votes Generation Speed (RTF) Organization License Training Data Supports Lyrics Access
1 riffusion-fuzz-1-0 1172.5 +99.1 / -62.7 144 5.60 Producer.ai Closed Unspecified True Proprietary
2 riffusion-fuzz-1-1 1087.3 +40.8 / -47.2 218 5.25 Producer.ai Closed Unspecified True Proprietary
3 preview-ocelot 1045.7 +75.9 / -82.9 90 5.42 Hidden Closed Unspecified True Proprietary
4 preview-jerboa 1034.4 +92.6 / -80.8 88 5.61 Hidden Closed Unspecified True Proprietary
5 acestep 660.1 +75.5 / -121.3 178 2.89 ACE Studio Apache 2.0 Unspecified True Open weights
Figure 2: Music Arena leaderboard (July 28 – Aug 31, 2025), plotting Arena Score (Y-axis) against Generation Speed (Median RTF, X-axis, log scale). Colors and shapes distinguish models by their training data and access type (open weights/proprietary), respectively. This visualization emphasizes the key tradeoff between model quality (score) and interactive latency (speed), an important consideration for creative music applications.
Music Arena was launched on July 28, 2025, and after collecting 1,051 user votes by the end of August, we released the first public leaderboard on September 19, 2025. Here, we provide an overview of the leaderboard’s structure, which is designed to address the evaluation challenges in the music domain. The detailed results of this initial leaderboard are presented in Table˜1, and the key tradeoffs are visualized in Figure˜2.
Public leaderboards for live evaluation platforms [12, 14, 16, 13] tend to contain similar attributes: an overall “Arena Score” (usually derived from the Bradley-Terry model [17]), the number of votes, the model provider, and the system license. There are a number of unique considerations in music that motivate presentation of additional attributes.
For LLM training, it can be broadly assumed that all providers are training on large-scale text data mined from the web. However, in music, we see significantly more diversity in training data across models. For example, some models are trained on licensed stock music [4] or publicly-available music under Creative Commons licenses [5], while some only specify the quantity (rather than the provenance) of their training data [21]. These differences induce an uneven playing field for training data, affecting not only a system’s performance in pairwise comparisons but also its standing within broader legal and ethical conversations. Accordingly, we include on our leaderboard a summary of available training data information for each model, including provenance and quantity.
In addition to training data, we include generation speed on our leaderboard. Tools that facilitate the creation of music are often designed to have low latency interactions. Accordingly, TTM providers may make different tradeoffs between quality and speed, depending on if they are targeting more consumption-oriented (prefer quality) or creative-oriented (prefer speed) applications. Speed is codified on our leaderboard by median real-time factor (RTF), where RTF is the ratio of seconds of music generated divided by seconds of wall clock time to generate. For example, a system that generates $30$ seconds of audio in $3$ seconds has an RTF of 1010x. Measuring speed via RTF gracefully handles variable length outputs, ensuring that systems are not punished for taking more time to generate longer audio.
The leaderboard interface allows users to view results segmented by model capability (e.g., “Instrumental” vs. “Vocal” tabs), allowing viewers to make principled decisions about models based on attributes of keen importance to their specific music application goals. Moreover, we emphasize these music-specific tradeoffs through visualization: a 22D scatter plot with speed on the X axis, Arena Score on the Y axis, and colors and shapes to emphasize training data and licensing information.
Appendix B Analysis of Initial Data Release
This section provides a more detailed look into user engagement patterns, prompt characteristics, and the types of music users create on the Music Arena platform.
B.1 User Engagement Distribution
Our analysis shows a long-tail distribution of user engagement. While many users contribute a small number of votes, a dedicated group of “power users” is responsible for a significant portion of the data. Table˜2 shows the distribution of votes submitted per user from our initial data release.
Table 2: Distribution of votes per user, based on 1,051 valid votes from 373 unique users.
Number of Votes Number of Users
1 193
2 72
3 44
4 24
5 8
6-10 18
11-20 10
21-50 4
B.2 Prompt Descriptiveness
Analyzing the 804 user-written prompts from valid, voted-on battles reveals they are typically concise. As shown in Table˜3, the raw data shows a median prompt length of 7 words, but the mean (18.68) is heavily skewed by a long tail of very descriptive prompts (max 1000 words). To get a more accurate picture of typical behavior, we removed 82 extreme outliers using the IQR method (threshold at 33 words). Table˜3 shows that the statistics for the remaining 722 prompts are much more focused, with a median length of 6 words and a mean of 8.27. Figure˜3 visualizes this post-filtered distribution, confirming that the typical Music Arena user prefers to express their creative ideas in a few words.
Table 3: Descriptive statistics for user prompt lengths (in words), before and after Interquartile Range (IQR) outlier removal.
Metric Raw Data (804 prompts) Post-IQR (722 prompts)
count 804 722
mean 18.68 8.27
std 54.82 6.87
min 1 1
50% (median) 7 6
max 1000 33
Figure 3: Distribution of user prompt lengths from voted battles (after removing outliers)
B.3 Commonly Requested Musical Concepts
By analyzing the keywords in 804 user-written prompts, we identified the most common musical genres, instruments, and moods requested by users. Table˜4 lists the most frequent keywords, and Figure 4 visualizes their prominence.
Figure 4: A word cloud of the most frequent keywords in user prompts from voted battles.
Table 4: Most frequent keywords in user-written prompts.
Keyword Frequency
bass 101
pop 98
vocals 81
piano 70
rock 69
dark 66
melodic 66
chorus 65
Appendix C Complete example of a Music Arena battle
@@
In Figure˜5 we show an example of a completed battle in Music Arena. Below we include the complete JSON log from our platform for that same battle, highlighting the detailed preference information that we collect. The corresponding audio is here: https://drive.google.com/drive/folders/1UlueXyaTmef2qw5zwdctXNVgKP9bFa1R?usp=sharing